# RF-3DGS: Wireless Channel Modeling with Radio Radiance Field and 3D Gaussian Splatting

### v1

@@ II-B Radiance Field Reconstruction Methods
where $C$ is the rendered pixel color, $c(t,d)$ is the view dependent color at position $t$, $d$ is the view direction, $\alpha(t)$ is the density, and $T(t)$ is the transmittance, representing the integration of density from the image plane to $t$. When rendering a solid object, the ray intersects the object as shown in the figure, causing the transmittance $T(t)$ to sharply drop from 1 to 0 due to the high density $\alpha(t)$ at the object’s surface. As a result, the portion of the ray behind the intersection has little contribution, $Con(t)=\alpha(t)T(t)$, to the pixel color. The free space portion before the intersection also contributes minimally due to the very small density $\alpha(t)$. This means that only the segment near the intersection significantly influences the pixel’s color, which aligns with our visual experience.
Figure 3: Principles of volume rendering.
However, directly modeling and integrating such continuous fields $c(t,d)$ and $\alpha(t)$ is infeasible for most volumetric data. In [41], the integral function is discretized into multiple segments of length $\delta_{i}$, and the discrete volume rendering is formulated as:
$C=\sum_{i=1}^{N}T_{i}\left(1-\exp(-\alpha_{i}\delta_{i})\right)c_{i},$ (3)
$T_{i}=\exp\left(-\sum_{j=1}^{i-1}\alpha_{j}\delta_{j}\right).$ (4)
To reconstruct the optical radiance field from images of the scene, we have a learnable representation capable of representing the radiance field and the geometry. This representation is subsequently optimized by iteratively rendering the rays corresponding to the images and performing gradient descent to minimize the loss between the rendered and ground truth colors. Furthermore, for more efficient training, a well-designed representation structure, rendering pipeline, and training strategies are essential.
@@
Beyond these techniques and representations, the fundamental philosophy behind how such methods capture scene geometry is straightforward: since a surface point on an object is visible from multiple views, the rays responsible for rendering this point converge from different camera positions to the same location. As a result, during gradient descent, only this surface point should receive an increase in density and an adjustment to its corresponding color, as this is the optimal solution to minimize the sum of the loss functions associated with those rays. This insight suggests that our radio spectra should also contain sufficient and accurate geometric information and maintain consistency across different views.
@@
II-C Array Signal Processing and Spatial Spectrum
Considering applications in the wireless communication domain, we assume the system employs a common wireless communication configuration, typically involving half-wavelength spaced Uniform Planar Arrays (UPA) operating with either a digital beamformer or a fully analog beamformer to generate baseband digital signals. Systems specifically designed for RF imaging/radar are not considered in this paper.
In a typical array signal processing scenario [42], there are $M$ signals incident from different directions onto the array, received by the $N$ elements of the UPA, and processed by the $L$ steering vectors, resulting in the desired signal intensities $\mathbf{y}(t)$ across the $L$ scanning directions. This process can be expressed as:
$\mathbf{y}(t)=\mathbf{W}^{H}\cdot\mathbf{A}(\mathbf{\Theta},\mathbf{\Phi})\cdot\mathbf{s}(t)+\mathbf{n}(t).$ (5)
The signal vector $\mathbf{s}(t)$ is defined as:
$\mathbf{s}(t)=\left[s_{1}(t),s_{2}(t),\dots,s_{M}(t)\right]^{T},$ (6)
where $s_{m}(t)$ represents the $m$-th incident signal characterized by its incident zenith angle $\theta_{m}$ and azimuth angle $\phi_{m}$. $\mathbf{n}(t)$ is the additive white Gaussian noise.
The matrix $\mathbf{A}(\mathbf{\Theta},\mathbf{\Phi})$ is a complex matrix of size $N\times M$, with each column corresponding to the array manifold vector of the $m$-th signal:
@@ II-C Array Signal Processing and Spatial Spectrum
Thus, the element responses are:
$\mathbf{x}(t)=\mathbf{A}(\mathbf{\Theta},\mathbf{\Phi})\cdot\mathbf{s}(t).$ (9)
The objective in our case is to find an appropriate steering vector set $\mathbf{W}^{H}=\left[w_{1},w_{2},\dots,w_{L}\right]$, that maximize the Signal-to-Interference-plus-Noise Ratio (SINR) of the signals in each corresponding AoA direction.
In our context of massive MIMO systems operating at mmwave bands within a fully scattering environment, we can obtain the estimated signal intensities $\mathbf{s}(t)$ across a set of scanning directions and then be organized into the desired spatial spectrum. However, directly using intensity as the pixel value introduces a notable issue. For EM wave propagation in free space, the pathloss follows the well-known Friis equation:
$P_{r}=P_{t}\frac{G_{t}G_{r}\lambda^{2}}{(4\pi d)^{2}},$ (10)
where $d$ is the distance between the Tx and Rx. This equation indicates that the received power decreases proportionally to $1/d^{2}$. Referring back to Fig. 3 and Eqn. 1, we note that along a ray, the rendered color is entirely independent of the distance between the object surface and the camera. This is consistent with our experience; for example, when viewing an LED bulb from 1 m and then 2 m away, the perceived brightness does not drop by four times.
This discrepancy between the Friis equation and volume rendering can be explained by the human eye’s response to light intensity, which follows a logarithmic-like curve and spans a dynamic range of approximately 100 dB [43]. Therefore, to adapt NeRF-style methods to our context, the spectrum must undergo a similar transformation. In this paper, we convert the AoA-gain spectrum to the dB scale to mimic such distance-independent decay.
II-D Spatial Spectrum
To accommodate different wireless systems, we explored various array signal processing techniques to provide an overview of spectra that different systems can generate, as shown in Fig. 4. The impact of these spectra on radio radiance field reconstruction will be further illustrated in the next section.
@@ II-D Spatial Spectrum
To address this issue, we tested the CBF spectrum tapered by a Hann window [42] and found that while side lobe levels were suppressed, interference remained problematic on the dB scale. Consequently, we adopted the NeRF2 configuration, using a linear scale transform and per-view normalization for both CBF and Tapered CBF (TCBF) spectra. While this configuration effectively represents AoA in small environments, it contradicts the distance-independent decay and view consistency assumptions of NeRF-style methods, resulting in poor performance in larger environments.
@@
where $\mathbf{R}=\mathbf{x}(t)\mathbf{x}^{H}(t)$ is the signal autocorrelation matrix. The optimal steering vector and the corresponding power spectrum can be written as:

### v2

@@ II-C Rendering the RRF Representation
where $T(t)$ is the transmittance, representing the accumulated density from the image plane to $t$, with $t_{n}$ and $t_{f}$ being the pre-defined lower and upper integration limits along the ray. Notably, as shown by the three curves in the lower-right part of Fig. 3, when the rendering ray intersects the blue ball surface with high density $\alpha(t)$, the transmittance $T(t)$ drops sharply from 1 to 0. As a result, the portion of the ray behind the intersection has little contribution, $Con(t)=\alpha(t)T(t)$, to the $\mathbf{c_{recv}}$. Similarly, the portion of the ray in free space before the intersection also contributes little due to their zero density. Thus, only the radio radiance of the first encountered surface point significantly influences the received Spatial-CSI, which aligns with our intuition.
II-D Reconstructing the RRF
To reconstruct the RRF $\mathcal{R}$ from a set of training radio spatial spectra captured at different array poses, we first need a learnable representation capable of encoding the geometry $\alpha(\mathbf{x})$ and the radio radiance $\mathbf{c}(\mathbf{x},\mathbf{d})$, while maintaining differentiability to the two functions. Additionally, the representation should be editable, concise, and easy to initialize.
Figure 4: Train a RRF representation from radio spatial spectra.
Now, assuming we have a well-initialized learnable RRF representation, this representation is iteratively optimized through gradient descent to minimize the loss between the rendered radio spatial spectra and the training spectra. While it is clear that the alpha-blending rendering equation is also differentiable to the two functions, allowing the loss and representation to be directly connected, the efficiency of this training process requires seamless cooperation between the representation structure, the rendering pipeline, and the training strategies. Thus, building on 3DGS, RF-3DGS addresses these challenges by utilizing explicit 3D Gaussian distributions and SH Functions to represent the RRF $\mathcal{R}$. For rendering, a parallelized rasterization pipeline is employed, along with adaptive training strategies modified specifically for RRF reconstruction.
@@ II-D Reconstructing the RRF
Beyond these techniques, the core philosophy behind how such methods capture the RRF $\mathcal{R}$ is straightforward. As illustrated in Fig. 4, since the red surface point on the object appears in multiple training radio spatial spectra, each spectrum contains a rendering ray that renders this point during optimization. As all these rays converge at the actual position of the surface point in 3D space, during gradient descent, only this surface point in the representation should receive an increase in density and its corresponding Spatial-CSI, as this is the optimal solution to minimize the loss functions associated with those rays. This insight suggests that our radio spatial spectra should contain sufficient and accurate geometric information while maintaining consistency across different views. However, wireless samples naturally lack geometric information, and such requirements necessitates expensive measurement systems and significant processing effort. RF-3DGS overcomes this challenge with a two-stage fusion training strategy to efficiently integrate geometric information from visual inputs into the RRF reconstruction.
Figure 5: RF-3DGS pipeline. (a) The extracted geometry representation from the visual data, consisting of millions of Gaussians, where the Gaussians are visualized by their contour surfaces. (b) The reconstructed radio radiance field, where the objects are illuminated by RF waves. (c) A demonstration of SH functions. (d) An example of the CSI-encoded spectrum and the corresponding decoded CSI spectra from the respective channels.
III Related Works
Ray tracing: Originally developed for computer graphics, ray tracing has long been adapted for radio propagation modeling [26, 27]. It models radio waves as rays to simulate complex RF interactions such as reflection, diffraction, and scattering [10, 11, 12]. Although ray tracing is suitable for a wide range of environments, it is limited by high computational complexity, stringent requirements for environmental information, and an inflexible pipeline that is difficult to calibrate.
NeRF2 reconstructs a “squared” radio radiance field that considers both “camera” and “light source” locations as inputs. However, its spectra, generated using CBF at 915 MHz with linear scale transformation and per-view normalization, suffer from strong interference, inconsistencies across views, lack of geometric information, and distance-dependent decay, all of which conflict with the requirements discussed earlier. To maintain effectiveness, NeRF2 requires a sample density 2000 times higher than in this work, negating the advantage of sparse sampling that NeRF offers. RayProNet [32] integrates environmental information as input and uses an explicit ”light probe” as the radio radiance field voxel. However, RayProNet is limited to producing only 2D path loss maps, without offering more detailed information.
IV Proposed RF-3DGS Framework
In the previous section (Sec. II-D), three critical challenges for efficient RRF reconstruction were identified. The first challenge is designing a learnable representation that can efficiently capture the geometry and the radio radiance of each object point. The second challenge involves developing an efficient rendering pipeline and training strategy to improve reconstruction speed and quality. The third challenge is the inherent lack of geometric information in radio spatial spectra. In the following subsections, we will introduce the representation structure, rendering pipeline, and training strategy of RF-3DGS, which are designed to address these three challenges.
IV-A 3D Gaussian Distributions
Unlike NeRF, which uses two neural networks to directly fit the density field function $\alpha(\mathbf{x})$ and the RRF function $\mathbf{c}(\mathbf{x},\mathbf{d})$ for all continuous points $\mathbf{x}$, 3DGS utilizes millions of learnable 3D Gaussian distributions (hereafter referred to as “Gaussians”) as the geometry representation and only reconstructs the radiance for each Gaussian center. In RF-3DGS, we adopt a similar configuration, using Gaussians as the basic primitives for the geometry. Each Gaussian has its own learnable parameters, including the basic density $\alpha_{g}$, the center (mean) of the Gaussian $\mathbf{x}$, and the shape of the Gaussian, which is typically defined by a $3\times 3$ covariance matrix $\boldsymbol{\Sigma}$. The density field of a single 3D Gaussian distribution can be represented as:
$\alpha(\mathbf{x})=\alpha_{g}\cdot e^{-\frac{1}{2}(\mathbf{x})^{\mathrm{T}}\boldsymbol{\Sigma}^{-1}(\mathbf{x})}$ (3)
However, due to the fact that the covariance matrix of a Gaussian must be positive semi-definite, it cannot be directly constrained during gradient descent. Therefore, the shape of the 3D Gaussians is represented by a more constrained set of parameters. The covariance matrix can be decomposed as:
$\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\mathrm{T}}\mathbf{R}^{\mathrm{T}}$ (4)
where $\mathbf{S}$ is a $3\times 3$ scaling diagonal matrix, and its diagonal elements consist of a 3D scaling vector $\mathbf{s}_{scal}$ that controls how the 3D Gaussian spreads along the three dimensions. The matrix $\mathbf{R}$ is a $3\times 3$ rotation matrix (an orthogonal matrix), which can be mapped with a quaternion $\mathbf{q}$ that controls the rotation of the Gaussian, as shown in Fig. 5. Both $\mathbf{s}_{scal}$ and $\mathbf{q}$ are easier to constrain within valid domains. For instance, large, anomalous Gaussians with excessive $\mathbf{s}_{scal}$ values can be excluded, and all $\mathbf{q}$ values can be normalized to unit length, ensuring validity.
Thus, during the training process, the parameters of the millions of Gaussians are optimized, “moving” and “transforming” these Gaussians to precisely represent the geometry of the environment. This is illustrated in Fig. 5(a), where the Gaussians are visualized by their contours of distribution density at 0.8 (the value of $e^{-\frac{1}{2}(\mathbf{x})^{\mathrm{T}}\boldsymbol{\Sigma}^{-1}(\mathbf{x})}$).
IV-B CSI-encoded Spherical Harmonics Function
Given the explicit object geometry consisting of Gaussians with exact positions, we only need to reconstruct the radio radiance from each Gaussian center. For this purpose, we use SH functions rather than a neural network. As mentioned earlier, the radio radiance $\mathbf{c_{x}}(\mathbf{d})$ at a point $\mathbf{x}$ is a spherical function, and such a complex spherical function can be adaptively approximated by SH function, which is a linear combination of basis SH functions. This representation is illustrated in Fig. 5(c). The number of basis functions is determined by the degree; the higher the degree, the more basis functions are available, allowing for the representation of more complex spherical functions. For our purpose, a degree of 3 is sufficient for RRF reconstruction.
In the Sec. II-C, we showed the rendering of single-value radio spatial spectra and mentioned that multi-modal Spatial-CSI requires more careful processing. This is mainly because, unlike the RGB values in normal images that already have the same distribution, the channel gain, delay, and AoD of an MPC have very different forms. First, the channel gain represents the ratio between transmitted power and received power for this MPC, and also indicates its importance. Therefore, the channel gain serves as the basis to encode the other CSI values so that our RRF representation can capture more details of the important MPCs.
However, directly using the absolute value of the channel gain as the pixel value introduces a significant issue. For EM wave propagation in free space, the channel gain (path loss) follows the well-known Friis equation:
$P_{r}=P_{t}\frac{G_{t}G_{r}\lambda^{2}}{(4\pi d)^{2}},$ (5)
where $d$ is the distance between the Tx and Rx. This equation indicates that the received power decreases proportionally to $1/d^{2}$. Referring back to Fig. 3 and Eq. (1), we note that along a ray, the rendered channel gain is independent of the distance between the object surface point and the array. This discrepancy arises because the radiance field method was originally designed to mimic visible light behavior. For visible light, when viewing an LED bulb from 1 meter and then 2 meters away, the perceived brightness does not drop by four times. This is explained by the human eye’s response to light intensity, which follows a logarithmic-like curve and spans a dynamic range of approximately 100 dB [33]. Therefore, the distance-dependent intensity decrease usually only results in a 1-2 dB drop in brightness, which is negligible. Consequently, in this paper, we convert the channel gain spectrum to the dB scale to mimic such distance-independent decay. Notably, this lossy transformation is not a compromise of the radiance field method’s limitations, but rather a way to take advantage of the visual intensity range. This transformation allows the spectrum to capture more details of the RRF, as shown in Fig. 8, where the dB scale spectrum more clearly describes the radio radiance field, at the cost of a few dB deviation in reconstruction.
Next, the dB-scaled channel gains are mapped to pixel values in the range [0, 255]. The upper limit of the channel gain is set to the free space propagation loss at 0.5 meters to avoid unexpected high values in testing and online operation. The lower limit is the lowest sampled value in the training samples, usually the noise level. For improved visualization, most demonstrations in this paper apply a “jet” colormap to single-channel radio spatial spectra.
Thus, along with the channel gain spectrum, the AoD and delay can be encoded using parallel channels. Specifically, AoD can be perfectly represented by our current RRF representation, as it only depends on the geometry of each ray. Therefore, we decompose it into two channels: one encoding the azimuth angle ($0^{\circ}$ to $360^{\circ}$) and the other encoding the zenith angle ($0^{\circ}$ to $180^{\circ}$).These twoangles of a MPC (pixel) then , which reflect the importance of this MPC.
For the delay channel, we currently define “delay” as the normalized delay of each MPC relative to the first-arrival MPC, ranging from 0 to 200 ns. MPCs exceeding 200 ns are discarded considering the indoor environment in this study. Although time of flight (ToF) would be preferable, it is difficult to acquire in field measurements, and it cannot be represented in the current RRF representation, as it is linearly related to the radio travel distance. The normalized delay alleviates this issue. For example, considering different Rx positions along a ray corresponding to the first-arrival MPC, the delay of later MPCs from other directions will slowly vary, which can be perfectly captured by the current RRF representation. However, once the different Rx positions are along a ray that does not correspond to the first-arrival MPC, the delays of other MPCs may deviate from the true delay. The larger the angle between the first-arrival MPC ray and the rays corresponding to later MPCs, the larger the deviations will be. However, such delay deviations are not significant, as shown in our later test of reconstructing delay spectra at unvisited locations.
Furthermore, the deviation in channel gain due to the dB scale and the deviation of delay representation can be further addressed with more in-depth development of the representation structure and rendering pipeline, which will be explored in future work.
IV-C Rendering RRF with the 3DGS Rasterization Pipeline
As the millions of 3D Gaussians collectively define a density field function, and the SH functions define the radio radiance of discrete object points (the means of the Gaussians), the alpha-blending rendering can be directly applied to the Gaussians along each ray, without relying on complex and lossy sampling strategies. However, as there are thousands of pixels in a spectrum, the rendering process must be parallelized to improve efficiency. Thus, instead of rendering the rays one by one which query all the Gaussians for each ray, each Gaussian is “splatted” onto the image plane, and all affected pixels are rendered simultaneously.
Specifically, for each individual Gaussian, the splatting is traditionally defined by the view transform matrix $\mathbf{W}$ and the perspective projection matrix $\mathbf{P}$. However, to simplify the calculation and increase stability, the Jacobian affine approximation $\mathbf{J}$ of the perspective projection is used to splat the 3D Gaussian onto the image plane, as given by:
$\mathbf{\Sigma^{\prime}}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W^{T}}\mathbf{J^{T}},$ (6)
where $\mathbf{\Sigma^{\prime}}$ is the covariance matrix of the projected 2D Gaussian. This affine approximation can be interpreted as each 3D Gaussian being first flattened to a 2D Gaussian by orthogonal projection along the camera depth axis, meaning that parts of the Gaussian do not appear larger as they get closer or smaller as they move further away, which is the cost of the affine approximation. Then, the 2D Gaussians are normally projected onto the 2D plane, following the perspective projection principle.
Next, an efficient rasterization rendering pipeline, well-suited for GPU computational structures, is employed. First, given the current rendering image plane and view frustum, the image plane is divided into $16~{}\times~{}16$ tiles, and the view frustum is divided accordingly. For each tile, only the Gaussians within the corresponding divided view frustum are selected (following a specific criterion) and sorted based on their depth to the plane. Finally, the computation within each tile consists of parallelized threads rendering the sorted Gaussians onto each pixel. Once a pixel is saturated with little transmittance, meaning that further Gaussians cannot affect this pixel, the thread rendering this pixel is terminated. Further details about the sorting process and GPU memory management for the millions of Gaussians can be found in [17].
IV-D Two-Stage Fusion Training
Given the rendered spectra from the learnable RRF representation and the training spectra, the loss function used in this work is a combined loss function. This function is defined as the $L1$ loss plus a differential Structural Similarity Index Measure (D-SSIM) term:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$ (7)
where $\lambda$ is a predefined parameter, set to 0.2 in this work. The gradient descent then backpropagates this loss through the differentiable rasterization pipeline to the learnable parameters of each Gaussian.
However, as mentioned earlier, a major challenge is that the radio spatial spectrum naturally lacks geometric information. From Fig. 8, even the ideal spectrum can only be considered a low-quality “photo”, with other spectra offering even less geometric detail. In RF-3DGS, we assume access to the radio spatial spectra and the visual data, which serve as a low-cost supplement to geometric information. The proposed two-stage fusion training process is designed to extract geometric information from the visual data to enhance the RRF representation.
Before the training begins, if the visual data consists of photos or video without camera intrinsic and extrinsic parameters, an initialization process is required to estimate these parameters and generate initial point clouds using SfM. Otherwise, this process can be skipped, allowing training to start from random point clouds.
Once initialization is complete, the first stage of training begins with a warm-up phase. During this phase, the Gaussians are trained using resolution-reduced images, enabling them to quickly capture global geometry rather than getting stuck on optimizing fine details. The image resolution is doubled twice, after 250 iterations and 500 iterations, until the original resolution is reached.
Following the warm-up, a periodic densification process refines the Gaussians every 100 iterations. In regions where the number of Gaussians is insufficient to represent intricate 3D geometry, the densification process splits large Gaussians and clones small Gaussians to create more learnable “ellipsoids” in those regions. Gaussians with very small density values or those with excessively large footprints in the views are removed to prevent the “floater” artifact problem, a common issue in radiance field reconstruction methods.
To manage the quantity of Gaussians, their basic density values $\alpha_{g}$ are reset towards zero every 3000 iterations. This reset allows key Gaussians to quickly regain their original density values during optimization, while abnormal or unnecessary Gaussians, whose density values increase slowly, are removed in subsequent densification steps.
After establishing a well-trained geometry representation, the second stage begins. In this stage, RF-3DGS uses the collected radio spatial spectra to train the CSI-encoded SH functions and the basic density of each Gaussian, while freezing the positions $\mathbf{x}$, rotation $\mathbf{q}$, and scaling $\mathbf{s_{scal}}$. This setup ensures that the geometry representation has the adaptive capability to handle inconsistencies between the visual and radio geometries (since certain materials may appear opaque in visual data but translucent under radio waves). After optimization, the new CSI-encoded SH functions can be interpreted as the objects being re-illuminated by the given Tx radio, with multiple channels corresponding to different Spatial-CSI properties.
The advantage of the proposed two-stage fusion training is not only that it requires low-cost visual data as a supplement, but also that it exhibits extrapolation capabilities, as shown in Fig. 6. This capability arises from the accurate visual geometric information, which provides the actual radiance source location, allowing vague or interfered radio spectra to be mapped to the correct locations rather than being assigned to vague floaters, as seen in NeRF2.
Figure 6: Extrapolation ability of the proposed two-stage fusion training. Both spectra are rendered using the same pinhole camera model. The ideal spectrum, serving as the training target, is limited by the maximum number of MPCs our RTX A6000 48 GB GPU can handle and the minimum array pattern lobe width needed for spectrum continuity.
@@ IV-D Two-Stage Fusion Training
IV-E Generating Radio Spatial Spectrum
Although RF-3DGS has largely addressed the challenges of RRF reconstruction, obtaining high-quality radio spatial spectra remains crucial. Considering applications in the wireless communication domain, we assume the system employs a common wireless communication configuration, typically involving half-wavelength spaced Uniform Planar Arrays (UPAs) operating with either a digital beamformer or a fully analog beamformer to generate baseband digital signals. Systems specifically designed for RF imaging/radar are not considered in this paper.
In this work, we constructed a digital twin framework to evaluate the effectiveness of RF-3DGS, comprising comprehensive field measurements in a 14 m $\times$ 15 m lobby and a digital replica of the same space with a full radio simulation pipeline, as shown in Fig. 7. The 60 GHz measurement data [34] was provided by the National Institute of Standards and Technology (NIST), USA. The simulator used is Sionna [35], an open-source communication simulator developed by NVIDIA, which is equipped with the Sionna RT [36] ray tracer and supports GPU acceleration.
As illustrated in Fig. 8, our simulation environment is a precise 3D replica of the NIST lobby with the same Tx position. First, the digital twin of the environment is reconstructed using photographs and LiDAR data. More specifically, each object is carefully segmented by material, and each material is assigned the corresponding EM properties. These EM models are then imported to the simulator, which uses ray tracing to simulate radio propagation paths. For the fixed Tx position, the radio propagation paths are captured at a set of sparse random Rx positions distributed throughout the lobby. Since the setup simulates a fully diffuse environment at 60 GHz, the number of paths between each Tx-Rx pair often exceeds 300,000, which limits a UPA array size to 8$\times$8. Next, the simulated MPCs, distributed in the angle domain (shown in Fig. 8), are received by the UPA. Finally, the received signals at each element are processed using various Array Signal Processing (ASP) techniques to generate the desired radio spatial spectrum. Further details on the simulation configuration and data processing are available in our open-source code.
Before introducing the ASP techniques, let’s provide a brief scenario description. In a typical ASP scenario [37], $M$ signals incident from different directions onto the array are received by the $N$ elements of the UPA. These signals are processed by $L$ steering vectors, resulting in the processed signals $\mathbf{y}(t)$ across the $L$ scanning directions. This process can be expressed as:
$\mathbf{y}(t)=\mathbf{W}^{H}\cdot\mathbf{A}(\mathbf{\Theta},\mathbf{\Phi})\cdot\mathbf{s_{signal}}(t)+\mathbf{n}(t).$ (8)
The signal vector $\mathbf{s_{signal}}(t)$ is defined as:
$\mathbf{s_{signal}}(t)=\left[s_{1}(t),s_{2}(t),\dots,s_{M}(t)\right]^{T},$ (9)
where $s_{m}(t)$ represents the $m$-th incident signal characterized by its incident zenith angle $\theta_{m}$ and azimuth angle $\phi_{m}$. $\mathbf{n}(t)$ is the additive white Gaussian noise. The matrix $\mathbf{A}(\mathbf{\Theta},\mathbf{\Phi})$ is a complex matrix of size $N\times M$, with each column corresponding to the array manifold vector of the $m$-th signal:
@@ IV-E Generating Radio Spatial Spectrum
The objective in our case is to find an appropriate steering vector set $\mathbf{W}^{H}=\left[\mathbf{w_{1}},\mathbf{w_{2}},\dots,\mathbf{w_{L}}\right]$, that maximize the Signal-to-Interference-plus-Noise Ratio (SINR) of the signals in each corresponding AoA direction. As we mentioned in Sec. II-C, we use a pinhole camera model to generate the radio spatial spectrum, which means we need to scan the AoAs of all the 200 $\times$ 300 pixels.
To accommodate different wireless systems, we explored various ASP techniques to provide an overview of spectra that different systems can generate, as shown in Fig. 8, more specific, the CBF, TCBF and MVDR patterns and spectra are simulated by a $8~{}\times~{}8$ UPA, the ideal spectra are generated by a synthesized ideal pattern with main lobe width 3, which is close to the pattern of a $64~{}\times~{}64$ UPA with CBF.
@@
To address this issue, we tested the CBF spectrum tapered by a Hann window [37] and found that while side lobe levels were suppressed, interference remained problematic on the dB scale.
@@
where $\mathbf{R}$ is the signal autocorrelation matrix. The optimal steering vector and the corresponding power spectrum can be written as: